What is the surface area of a cube with a 4 foot diagonal
Answer by nyc_function(2741)

Use the formula:
Surface Area of a Cube = 6 a^2
The diagonal is the hypotenuse of a right triangle in the cube.
The legs are represented by the letter a.
a^2 + a^2 = 4^2
2a^2 = 16
a^2 = 16/2
a^2 = 8
sqrt{a^2} = sqrt{8}
a = 2(sqrt{2})
Now that we know the value of a, plug it into the formula for the surface area of a cube.
Surface Area of a Cube = 6 (2(sqrt{2})^2
Can you finish now?